OpenStudy (anonymous):

George is building a kite. He has a rectangular piece of cloth that is 6 inches wide and 11 inches long. He needs to place a stick on the diagonal of the rectangle to make his kite. How long of a stick does he need? Round to the nearest tenth of an inch. I got 12.5 as my answer. Is that correct? ^^

OpenStudy (anonymous):

6^2 + 11^2 = 36 + 121 = 157 sqrt(157) is about 12.52

OpenStudy (anonymous):

yea that's correct

OpenStudy (anonymous):

Thank you! :D